OpenStudy (anonymous):

Anna mixes the letters S, E, L, E, C, T, E, and D thoroughly. Without looking, Mary draws one letter. Expressed as a fraction, decimal, and percentage, what is the probability that E will not be the letter Mary selects? A 3/8 0.375, 37.5% B 8/5, 1.6, 16% C 5/8, 0.625, 62.5% D 8/3, 2.667, 26.667%

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

number of letters that are not E = 5 total number of letters = 8 5/8 = 0.625 = 62.5%

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

i have one last question

OpenStudy (anonymous):

no problem :)....ask away

OpenStudy (anonymous):

Tammy mixes the letters S, C, H, O, O, and L thoroughly. Without looking, Allen draws one letter. Expressed as a fraction, decimal, and percentage, what is the probability that Allen will not select a consonant? (1 point) A.5/6, 0.833, 83.3% B.2/3, 0.667, 66.7% C.1/3, 0.333, 33.3% D.3/2, 1.5, 15%

OpenStudy (anonymous):

number of letters that are not consonants = 2 total of letters = 6 2/6 = 1/3 = 0.333 = 33.3%

OpenStudy (anonymous):

can you help me out with some other questions

OpenStudy (anonymous):

if so i will give you as many medals as i can plz pretty plz

OpenStudy (anonymous):

I can try....I am not the best at the more complicated probabiliyies

OpenStudy (anonymous):

typo..probabilities

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you, here they come

OpenStudy (anonymous):

Suppose you have a wallet full of money. If the probability of picking a $10 bill is 4/15 and the probability of picking a $5 bill is 1/3 what is the probability of picking a $20 bill A 1/15 B 1/5 C 4/15 D 2/5

OpenStudy (anonymous):

I do not know this one...sorry :(

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

next one

OpenStudy (anonymous):

I am not the best at probabilities...only the easier ones

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you might want to try to get somebody over here that knows more about this stuff

OpenStudy (anonymous):

A quality supervisor for a factory that makes video games examines a sample of 82 video games. He finds that 7 video games are scratched. Which is the best prediction of the number of video games that will be scratched in a shipment of 6,000? A 500 B 512 C 556 D 610

OpenStudy (anonymous):

A yogurt shop offers 2 different flavors of yogurt and 11 different toppings. How many choices are there for a single serving of yogurt with one topping A 20 B 11 C 13 D 22

OpenStudy (anonymous):

7/82 = x/6000 cross multiply 82x = (6000)(7) 82x = 42000 x = 42000/82 x = 512 (answer B) ============== I think you just multiply 2 x 11 = 22.....I think

OpenStudy (anonymous):

Eric has a pair of identical number cubes. The faces of each cube are numbered 1–6. Eric will roll the cubes both at the same time. What is the probability that the numbers showing face-up after the roll will be multiples of 3 A 2/9 B 1/3 C 1/6 D 1/9

OpenStudy (anonymous):

And thanks one the other one

OpenStudy (anonymous):

2/6 * 2/6 = 4/24 which reduces to 1/6

OpenStudy (anonymous):

omg your good thank you, and just a few more

OpenStudy (anonymous):

lets see what you got :)

OpenStudy (anonymous):

How many different ways can you and five friends sit in your assigned seats when you go to a concert A 6!:120 B 6!: 720 C 5!: 120 D 5!: 30

OpenStudy (anonymous):

you lost me on this one...sorry :(

OpenStudy (anonymous):

I want to say it is either A or B, only because there is 6 people involved...but that is just a guess

OpenStudy (anonymous):

im saying my lost too lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok next one lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

How many different ways can a band teacher select the lead and co-lead trumpet player from a group of 12 trumpet players? A 132 B 264 C 312 D 448

OpenStudy (anonymous):

When building a house, you have the choice of seven different floor plans, brick or stone exterior, and carpet or hardwood flooring in the house. How many choices are there for one house A 7 B 28 C 4 D 14

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

ok...this is going over my head....your gonna need someone else to help you because if I try to help, it is gonna be like " the blind leading the blind ". I am so sorry.

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

but thanks for your help

OpenStudy (anonymous):

let me see if i have one last question

OpenStudy (anonymous):

I wish you the best of luck with these. Good - luck

OpenStudy (anonymous):

ok....I can look at it...can't make no promises

OpenStudy (anonymous):

A bag contains 2 white marbles and 7 purple marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. a. What is the probability of selecting a purple marble and then a white marble? b. What is the probability of selecting two white marbles? c. Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row? Show your work

OpenStudy (anonymous):

might be kind of easy

OpenStudy (anonymous):

well...I know the chances of drawing a purple marble on the first try is 7/9.....and the chance of drawing a white marble on the second try is (I think) 2/8 or 1/4. So I think you just multiply them. 7/9 * 1/4 = 7/36....you will need a second opinion. chance of drawing 2 white marbles.....2/9 * 2/8 = 4/72 = 1/18 as for the last part...I don't know about showing the work...but I believe there is a greater chance of drawing 2 purple marbles because there is more purple marbles then there is white marbles. I am not 100% on these, it would be wise of you to get a second opinion.

OpenStudy (anonymous):

Good-luck with these :)

OpenStudy (anonymous):

thank you so much and thats all